About 3.6 babies are born each day at a local hospital. Assuming the Poisson distribution, what is the probability that no babies are born today?

Round to 4 decimal places.
Answer #1

In a local hospital ,3.6 babies are born each day.

X: Number of babies born in a day.

X ~ Poisson(λ-3.6)

The probability mass function of Poisson distribution is

) ; 0.1.2.3.

Now , the probability that no babies are born today is

-3.6 3.60 -e-3.6 = 0.0273 * P(X = 0)--еч.t) * 3.60